16 & 17
016 (part 1 of 2) 10.0 points Objects with masses of 159 kg and 476 kg are separated by 0.326 m. A 27.9 kg mass is placed midway between them 476 kg 159 kg 27.9 kg 0.326 m Find the magnitude of the net gravitational force exerted by the two larger masses on the 27.9 kg mass. The value of the universal gravi- tational constant is 6.672 x 10 11 N. m2kg2 Answer in units of N. 017 (part 2 of 2) 10.0 points Leaving the distance between the 159 kg and the 476 kg masses fixed, at what distance from the 476 kg mass (other than infinitely remote
Answer #1

16) for the net force on the 27.9 Kg mass

Force = force due to 476 Kg - force due to 159 Kg

Force = 6.67 *10^-11 * 27.9 * (476/(0.326/2)^2 - 159/(0.326/2)^2)

Force = 2.22 *10^-5 N

the magnitude of force acting on the middle is 2.22 *10^-5 N

17)

let the distance is x a

for the force on the third mass to be zero

force due to 159 Kg = force due to 476 Kg

G * 159 * m/(0.326 - x)^2 = G * 476 * m/x^2

159/(0.326 - x)^2 = 476/x^2

solving for x

x = 0.206 m

the distance from the 476 Kg mass is 0.206 m